If the De Broglie wavelength of an electron is equal to 400 nm calculate the velocity of the electron. Assume that the electron's speed is non-relativistic.
Answer: 1832.42 m/s

If the kinetic energy of an electron is 400 eV, calculate its De Broglie wavelength. For this non-relativistic electron you must first calculate its velocity from the general kinetic energy equation. Then you can find the De Broglie wavelength of the electron.

I cannot figure out the second part, please explain!
Answer #1


part B : use the Formua for KInetic Energy KE = P^2/2m

where p is momentum also P = h/L

where h is plancks consant

L is wavelength

so

KE = h^2/2m L^2

L^2 = h^2/2mKE


waveength L = h/sqrt(2mKE)

L = 6.626 e-34/(sqrt(2 * 9.11 e -31 * 400 * 1.6 e -19)

L = 6.13 e -11 m